OpenStudy (anonymous):

what is the value of x in the equation √3 +x-5 = -2

OpenStudy (anonymous):

\[\sqrt 3+x-5 = -2 \rightarrow \sqrt 3+x = 5-2 \rightarrow \sqrt 3+x = 3\] \[\rightarrow x = 3 -\sqrt 3 = 3- 1.732 =1.268\] @Buchanan

OpenStudy (anonymous):

my choices are a 3 b 46 c 6 d 12

OpenStudy (anonymous):

Pls check your question again and then post

OpenStudy (anonymous):

Is your question written like \[\sqrt{3+x} -5=-2\]

OpenStudy (anonymous):

radical 3 +5 -5 = -2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

\[\sqrt{3+x} -5=-2 \rightarrow \sqrt{3+x} =5-2 \rightarrow \sqrt{3+x} =3\] Squaring both sides we find \[( \sqrt{3+x})^2 =(3)^2 \rightarrow 3+x = 9 \rightarrow x= 9-3 \rightarrow x= 6\] Hence x=6 is the required answer. @Buchanan

OpenStudy (anonymous):

thank you